Oxalic acid, H2C2O4 has a pKa1 of 1.25 and a pKa2 of 4.14.

When sodium hydrogen oxalate is put into water, will the solution be acidic, basic, or neutral? Justify your answer with some equations and calculations and words.
Answer #1

For diprotic acid like Oxalic acid, H2C2O4

H2C2O4 H+ + HC2O4-   pKa1 = 1.25

HC2O4- H+ + C2O42-    pKa2 = 4.14

Sodium hydrogen oxalate, NaHC2O4 sits between H2C2O4 and Na2C2O4 in a titration curve of oxalic acid and sodium hydroxide.

Hence, the pH of NaHC2O4 is the between pH = pKa1 and pH = pKa2.

Thus, the pH of sodium hydrogen oxalate = (pKa1 + pKa2)/2 = (1.25 + 4.14)/2 = 2.70

Therefore, when sodium hydrogen oxalate is put into water, will the solution be acidic as pH of the solution will be 2.70 which is less than pH 7.